OpenStudy (anonymous):

x^3+2x^2=16x+32. solve polynomial equation by factoring then using zero product principle. what is the solution set?

OpenStudy (anonymous):

\[x ^{3}+2x ^{2}=16x+32\]

OpenStudy (radar):

\[x ^{2}(x+2)=16(x+2)\]divide both sides by (x+2) getting: \[x ^{2}=16\] \[x=\pm 4\]

OpenStudy (radar):

I believe they work: -64+32=-64+32 64+32=64+32

OpenStudy (radar):

Did you follow that lar0213

OpenStudy (radar):

I didn't show an intermediate step \[x ^{2}=16\]\[\sqrt{x ^{2}}=\sqrt{16}\]\[x=\pm4\] That should make it more understandable. Good luck with your studies.